OpenStudy (anonymous):

PLZ HELPP! What is the equation of the line in slope-intercept form? the line perpendicular to y = 1/3x + 5 through (2, 1) A. y = 1/3x + 7 B. y = -3x + 7 C. y = 3x + 7 D. y = -1/3x + 7

jimthompson5910 (jim_thompson5910):

the slope of y = 1/3x + 5 is what?

OpenStudy (anonymous):

-3?

jimthompson5910 (jim_thompson5910):

no it's just the number in front of the x

OpenStudy (anonymous):

just 3

jimthompson5910 (jim_thompson5910):

more like 1/3

OpenStudy (anonymous):

im sorry im just so lost havent done this in yearss

jimthompson5910 (jim_thompson5910):

the slope of y = 1/3x + 5 is 1/3

jimthompson5910 (jim_thompson5910):

now you flip the fraction to get 3/1 then you flip the sign to get -3/1 = -3

OpenStudy (anonymous):

so it would be A then correct?

jimthompson5910 (jim_thompson5910):

so the original slope is 1/3 the perpendicular slope is -3

jimthompson5910 (jim_thompson5910):

there's only one answer choice that has a slope of -3

OpenStudy (anonymous):

oo gotcha thank you

jimthompson5910 (jim_thompson5910):

you're welcome